Diffusion Morning Report 10/27/11

Ted Berg tberg at stanford.edu
Thu Oct 27 06:31:23 PDT 2011


Good Morning all,
     Looking good this morning. TEL is down for water prep work. Tylan 
Poly is down for MFS and pressure alarms. Epi had hazy wafers. MST 26 
sensor repaired-GeH4 to LPs back on.
     Leftovers:J is working on NH3 to fijis. No other major issues at 
this time.

					
	Date: 	10/27/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*up* 	** 	*10.49/30* 	** 	**
Tylan 8-SiGe 	*up* 	** 	*3.76/30* 	** 	**
Tylan 9-Poly 	*down* 	*MFS-pressure errors* 	*3.87/30* 	** 	*1*
Tylan 10-Nitride 	*yellow* 	** 	*2.04/30* 	** 	**
Tylan 11-LTO 	*yellow* 	** 	*10.25/30* 	** 	**
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*up* 	** 	** 	** 	**
Thermco 3 	*down* 	** 	** 	** 	**
Thermco4 	*down* 	** 	** 	** 	**
Thermco Poly1 	*yellow* 	** 	** 	*
* 	**
Thermco Poly2 	*yellow* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi 	*yellow* 	*lla issues,Hazy wafers* 	** 	** 	**
TEL Spa 	*down* 	*Water Work* 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
SEM 4160 	*up* 	** 	** 	** 	**
ZYGO 	*up* 	** 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*B on line pressure 733 psi* 	*changeover working 
well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20111027/138f8fcf/attachment.html>


More information about the diffusion mailing list